A sample of 81 observations has a mean 23 and standard deviation                                

 3.6, Test the Null Hypothesis H0 : μ = 25 against the alternative

 H1:  μ < 25, assuming α = .01


1
Expert's answer
2021-11-01T16:10:31-0400

"n=81 \\

\bar{x} = 23 \\

s = 3.6 \\

H_0: \mu = 25 \\

H_1: \mu < 25"

Test-statistic

"t= \frac{\bar{x} - \mu}{s / \sqrt{n}} \\

t = \frac{23-25}{3.6 / \sqrt{81} } \\

= \frac{-2}{0.4333} \\

= -1.566 \\

α=0.01 \\

df=n-1=81-1=80"

One-tailed test

"t_{crit} = 2.373"

Reject the null hypothesis if "t \u2264 -t_{crit}"

"t= -1.566 > t_{crit} = -2.373"

We accept the null hypothesis at 0.01 level of significance.